V=455.9
What is the radius and diameter of the basketball?
Thank you,
Clewis

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

Volume of a sphere = (4/3)(pi)r^3
.
The problem told us that the volume = 455.9
.
455.9 = (4/3)(3.14)r^3
1367.7 = (4)(3.14)r^3
341.925 = (3.14)r^3
341.925 = (3.14)r^3
108.893 = r^3
4.78 inches = r (radius)
.
Diameter = 2r = 2(4.78) = 9.55 inches (diameter)

Answer by Alan3354(69443)

V=455.9
What is the radius and diameter of the basketball?
-----------
Assuming the ball is a sphere, the volume of a sphere is (4/3)PIr^3
455.9 = 4*PI*r^3/3
r^3 = 455.9*3/(4*PI)
r^3 = 108.838
r = 4.77 units
d = 9.54 units